# Color Equivariant Network

### v1

@@ 1 Introduction
The tremendous progress of image classification in the last decade can be readily attributed to the development of deep convolutional neural networks [1; 2; 3]. The highly nonlinear mapping and large parameter space does not lend itself to interpretation easily, but even in early networks, representations of geometry and color were observed and recognized for their importance in representation learning [1; 4]. While a relatively large body of literature has worked to improve the robustness of convolutional neural networks to geometric transformations [5; 6; 7; 8; 9; 10], improving their robustness to perceptual variation has garnered considerably less attention.
@@ 2 Related Work
The strong generalization performance of convolutional neural networks on image processing tasks is attributed, in part, to the equivariance of planar convolution to 2D translations. This insight has garnered considerable attention and led to a strong interest in the development of convolutional neural networks equivariant to other symmetry groups [16; 17]. Previous works introduce group equivariant convolutional networks for finite and continuous groups. A framework for finite group convolution was developed in [6], and demonstrated for the 2D rotation, and reflection groups. The authors achieve equivariance to specific symmetry groups by convolving input images and feature maps with the group orbit of a learned filter bank. In [18], the authors design for equivariance to the continuous 2D rotation group by constraining filter representations to have circular harmonic structure.
@@
Other works encourage transformation equivariance using a soft penalty term on representation dissimilarity [4; 27]. These methods are separate from the group equivariant network literature, and different from what we propose. Our model leverages the geometric structure of hue and saturation to design a hue-saturation equivariant network bringing the benefits of group equivariant networks to perceptual transformations.
@@
Several prior works attempt to mitigate the effect of color variation in image processing tasks, for example, in [28], the authors represent images in a color space where pixel values are invariant to changes in luminance for luminance invariant image segmentation, and in [14], the authors penalize changes in their latent representations due to color variation for color invariant skin lesion identification. While color invariance resolves the effect of color variation, color has been identified as an important cue in neural network representation learning [12; 13] which is discordant with the goal of color invariance. In contrast to these approaches, our work leverages the geometric structure of hue, and saturation to construct a convolutional neural network equivariant to variations in these quantities by design.
@@ 3 Preliminaries
The equivariance of our network to hue shifts is illustrated in Figure 1(a).
@@ 4 Method
[…]
(a) Hue equivariance
(b) Lifting layer
Figure 1: Color equivariant network. (a) The equivariance of our hue equivariant model is illustrated by the commutativity of the (hue) rotation and neural network mapping. A hue rotation of 90 in the input image space (top-left to bottom-left), results in a feature map rotation at each layer of the network (top-right to bottom-right). (b) An input image (left) is “lifted” to the hue-saturation group (right) by shifting its hue and saturation values.
@@
An element of the hue group acts on a function on the discrete hue group by the group action $\phi_{h}:H_{N}\times Y\rightarrow Y$, which “rotates" the function on the group. Concretely, for a function $f$ on the discrete hue group $H_{N}$ defined as the concatenation of functions $f=(f_{1},\dots,f_{N})$, the action of an element $h_{i}$ in the hue group $H_{N}$ is given by
@@
An element of the saturation group acts on a function on the discrete saturation group by the group action $\phi_{s}:S_{N}\times Y\rightarrow Y$, which “translates" the function on the group. Concretely, for a function $f$ on the discrete saturation group $S_{N}$ defined as the concatenation of functions $f=(f_{1},\dots,f_{N})$, the action of an element $s_{i}$ in the saturation group $S_{N}$ is given by
@@
An element of the hue-saturation group acts on a function on the discrete saturation group by the group action $\phi_{hs}:H_{N}\times S_{M}\times Y\rightarrow Y$, which “rotates" and “translates" the function on the group. Concretely, for a function $f$ on the discrete hue-saturation group $H_{N}\times S_{M}$ defined as the concatenation of functions $f=(f_{11},\dots,f_{1M},f_{21},\dots,f_{NM})$, the action of an element $(h_{i},s_{j})$ in the hue-saturation group $H_{N}\times S_{M}$ is given by
@@
The first layer of a group convolutional neural network “lifts" the input image to the group (see Figure 1(b)). We can lift an input image to the product space of the image grid $\mathbb{Z}^{2}$, and discretized hue shifts $H_{N}=\{h_{0},h_{1},\dots,h_{N}\}$, by convolving with hue shifted filters,
@@ 5.1 Hue-shift MNIST: Global hue equivariance
(a) Z2CNN    (b) Hue-CNN    (c) Hybrid
Figure 3: Comparison of feature maps for equivariant and non-equvariant architectures. In the conventional network (a), the relationship between feature maps is unclear. In our model (b), a hue shift in the input results in a predictable hue shift of the internal representations. In the hybrid network (c), the feature maps appear approximately invariant to hue shift.
5.2 Hue-shift 3D Shapes: Local hue equivariance
@@ 5.3 Camelyon17: Saturation shift in the wild
We present experimental results for four discretizations of the saturation space. Our choice of discretizations, i.e., $d\in\{\nicefrac{{1}}{{20}},\nicefrac{{1}}{{10}},\nicefrac{{3}}{{20}},\nicefrac{{1}}{{2}}\}$, are determined with consideration of the saturation range of the training set. For a given choice of discretization $d$, an element $s_{i}\in S_{N}$ shifts the saturation of an input image by $s_{i}=i*d*255$. We limit computational expense of lifting and convolution by performing these operations over the truncated set $\{s_{-1},s_{0},s_{1}\}$.
ResNet50 Sat-$d_{1}$* Sat-$d_{2}$* Sat-$d_{3}$* Sat-$d_{4}$*
Error 28.91 (7.58) 19.60 (2.08) 16.08 (2.68) 24.53 (7.13) 23.57 (6.04)
Params 23.5M 23.3M 23.3M 23.3M 23.3M
Model performance is reported in Table 3. The performance of our saturation-equivariant models out perform all other models, and the performance of our hue-equivariant model is on par with the baseline model.
@@ 5.4 CIFAR-10: Color shift in the wild
We evaluate the performance of our model in the presence of natural hue shifts on the CIFAR-10 [34] classification dataset. The dataset consists of 60k examples, 40k of which are used for training, 10k for validation, and 10k for testing.
ResNet44 Hue-4* Sat-$d_{1}$* Sat-$d_{2}$* Sat-$d_{3}$*
Error 7.86 (1.14) 8.83 (0.64) 10.45 (1.76) 10.26 (0.43) 9.24 (0.27)
Params 2.64M 2.63M 2.55 2.55 2.55M
Hue-Sat-$d_{1}$* Hue-Sat-$d_{2}$* Hue-Sat-$d_{3}$* Hue-3* CEConv
Error 10.68 (0.78) 11.13 (0.84) 11.16 (0.38) 8.42 (0.39) 8.86 (0.33)
Params 2.50M 2.50M 2.50M 2.55M 1.15M
Figure 4: CIFAR-10 automobile class sorted by hue.
Model performance is reported in Table 4. The performance of our color-equivariant models are comparable to the baseline model.
@@ 6.1 Limitations and future work
While our method is equivariant to hue and saturation by design, it is not guaranteed to be equivariant to luminance. However, since luminance is computed as a linear combination of the red, green and blue channel values of an input image, if the preprocessing statistics (e.g., data mean) on the training set are valid for the testing set, luminance is approximately unchanced.
@@ A.1 Hue-shift MNIST
We introduce the Hue-shift MNIST dataset to evaluate the performance of our equivariant architectures in the presence of global hue shift. We show select examples from the in-distribution and out-of distribution datasets in Figures 5(a) and 5(b).
@@ A.2 Hue-shift 3D Shapes
We introduce the Hue-shift 3D Shapes dataset to evaluate the performance of our equivariant architectures in the presence of local hue shift. We show select examples from the in-distribution and out-of distribution datasets in Figures 6(a)- 6(c).
@@ B.1 Hue-shift MNIST
We train our equivariant networks and the conventional architectures for 5 epochs with a batch size of 128. We optimize over a cross-entropy loss using the Adam optimizer Kingma and Ba [2014] with $\beta_{1}=0.9$, and $\beta_{2}=0.999$. We use an initial learning rate of $10^{-3}$ for the $\mathbb{Z}^{2}$ network, and $10^{-4}$ for the hue-equivariant network. We train CEConv architectures using the hyperparameters reported in Lengyel et al. [2024].
@@ B.2 Hue-shift 3D Shapes: Local hue equivariance
For this task we compare against the ResNet18 architecture proposed in He et al. [2016], and the Z2CNN and ResNet44 architectures proposed in Cohen and Welling [2016]. Our hue-equivariant architectures are designed with the same network structure as the baselines, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our hue-equivariant architectures, we perform hue group pooling to yield an hue-invariant representation.
Following Wiles et al. [2021], we train the ResNet architectures for 100k iterations with a batch size of 128. We optimize over a cross-entropy loss using SGD with a learning rate of $10^{-2}$. Images are down-sampled by a factor of 2 to train the CNN architectures. CNN architectures are trained as described in Section B.1.
@@ B.3 CIFAR-10: Color shift in the wild
For this task, we compare against the ResNet44 architecture proposed in Cohen and Welling [2016]. Our hue-, saturation- and hue-saturation-equivariant architectures have the same number of layers as ResNet44, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our networks, we perform group pooling to yield group invariant representation.
We train our equivariant networks and the conventional architectures for 300 epochs, and a batch size of 128. We optimize over a cross-entropy loss using SGD with an initial learning rate of $10^{-1}$ and a cosine-annealing scheduler. We train CEConv architectures using the hyperparameters reported in Lengyel et al. [2024].
@@ B.4 Camelyon17: Saturation shift in the wild
For this task, we compare against the ResNet50 architecture He et al. [2016]. Our hue- and saturation-equivariant architectures have the same number of layers as ResNet50, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our hue- and saturation-equivariant architectures, we perform a group pooling to yield group invariant representation.
We train our equivariant networks and the conventional architectures for 10k iterations with a batch size of 32. We optimize over a cross-entropy loss using the Adam optimizer Kingma and Ba [2014] with an initial learning rate of $10^{-2}$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$. We train CEConv architectures using the hyperparameters reported in Lengyel et al. [2024].

### v2

@@ 1 Introduction
[…]
Figure 1: Hue equivariance. The equivariance of our hue equivariant model is illustrated by the commutativity of the (hue) rotation and neural network mapping. A hue rotation of 90 in the input image space (top-left to bottom-left), results in a feature map rotation at each layer of the network (top-right to bottom-right).
The tremendous progress of image classification in the last decade can be readily attributed to the development of deep convolutional neural networks [24, 30, 19]. The highly nonlinear mapping and large parameter space does not lend itself to interpretation easily, but even in early networks, representations of geometry and color were observed and recognized for their importance in representation learning [24, 26]. While a relatively large body of literature has worked to improve the robustness of convolutional neural networks to geometric transformations [4, 7, 16, 28, 20, 17], improving their robustness to perceptual variation has garnered considerably less attention.
@@ 2 Related work
The strong generalization performance of convolutional neural networks on image processing tasks is attributed, in part, to the equivariance of planar convolution to 2D translations. This insight has garnered considerable attention and led to a strong interest in the development of convolutional neural networks equivariant to other symmetry groups [22, 9]. Previous works introduce group equivariant convolutional networks for finite and continuous groups. A framework for finite group convolution was developed in [7], and demonstrated for the 2D rotation, and reflection groups. The authors achieve equivariance to specific symmetry groups by convolving input images and feature maps with the group orbit of a learned filter bank. In [36], the authors design for equivariance to the continuous 2D rotation group by constraining filter representations to have circular harmonic structure.
@@
Other works encourage transformation equivariance using a soft penalty term on representation dissimilarity [26, 18]. These methods are separate from the group equivariant network literature, and different from what we propose. Our model leverages the geometric structure of hue and saturation to design a hue-saturation equivariant network bringing the benefits of group equivariant networks to perceptual transformations.
@@
Several prior works attempt to mitigate the effect of color variation in image processing tasks, for example, in [6], the authors represent images in a color space where pixel values are invariant to changes in luminance for luminance invariant image segmentation, and in [27], the authors penalize changes in their latent representations due to color variation for color invariant skin lesion identification. While color invariance resolves the effect of color variation, color has been identified as an important cue in neural network representation learning [11, 10] which is discordant with the goal of color invariance. In contrast to these approaches, our work leverages the geometric structure of hue, and saturation to construct a convolutional neural network equivariant to variations in these quantities by design.
@@ 3 Preliminaries
The equivariance of our network to hue shifts is illustrated in Figure 1.
@@ 4 Method
Figure 2: Lifting layer. An input image (left) is “lifted” to the hue-saturation group (right) by shifting its hue and saturation values.
@@
An element of the hue group acts on a function on the discrete hue group by the group action $\phi_{h}:H_{N}\times Y\rightarrow Y$, which “rotates” the function on the group. Concretely, for a function $f$ on the discrete hue group $H_{N}$ defined as the concatenation of functions $f=(f_{1},\dots,f_{N})$, the action of an element $h_{i}$ in the hue group $H_{N}$ is given by
@@
An element of the saturation group acts on a function on the discrete saturation group by the group action $\phi_{s}:S_{N}\times Y\rightarrow Y$, which “translates” the function on the group. Concretely, for a function $f$ on the discrete saturation group $S_{N}$ defined as the concatenation of functions $f=(f_{1},\dots,f_{N})$, the action of an element $s_{i}$ in the saturation group $S_{N}$ is given by
@@
An element of the hue-saturation group acts on a function on the discrete saturation group by the group action $\phi_{hs}:H_{N}\times S_{M}\times Y\rightarrow Y$, which “rotates” and “translates” the function on the group. Concretely, for a function $f$ on the discrete hue-saturation group $H_{N}\times S_{M}$ defined as the concatenation of functions $f=(f_{11},\dots,f_{1M},f_{21},\dots,f_{NM})$, the action of an element $(h_{i},s_{j})$ in the hue-saturation group $H_{N}\times S_{M}$ is given by
@@
The first layer of a group convolutional neural network “lifts” the input image to the group (see Figure 2). We can lift an input image to the product space of the image grid $\mathbb{Z}^{2}$, and discretized hue shifts $H_{N}=\{h_{0},h_{1},\dots,h_{N}\}$, by convolving with hue shifted filters,
@@ 5 Experiments
(a) Z2CNN      (b) Hue-CNN      (c) Hybrid
Figure 4: Comparison of feature maps for equivariant and non-equivariant architectures. In the conventional network (a), the relationship between feature maps is unclear. In our model (b), a hue shift in the input results in a predictable hue shift of the internal representations. In the hybrid network (c), the feature maps appear approximately invariant to hue shift.
@@ 5.1 Hue-shift MNIST: Global hue equivariance
5.2 Hue-shift 3D Shapes: Local hue equivariance
Error Params
ResNet50 28.91 (7.58) 23.5M
Sat-$d_{1}$* 19.60 (2.08) 23.3M
Sat-$d_{2}$* 16.08 (2.68) 23.3M
Sat-$d_{3}$* 24.53 (7.13) 23.3M
Sat-$d_{4}$* 23.57 (6.04) 23.3M
Figure 5: CIFAR-10 automobile class sorted by hue.
@@ 5.3 Camelyon17: Saturation shift in the wild
We present experimental results for four discretizations of the saturation space. Our choice of discretizations, i.e., $d\in\{\nicefrac{{1}}{{20}},\nicefrac{{1}}{{10}},\nicefrac{{3}}{{20}},\nicefrac{{1}}{{2}}\}$, are determined with consideration of the saturation range of the training set. For a given choice of discretization $d$, an element $s_{i}\in S_{N}$ shifts the saturation of an input image by $s_{i}=i*d*255$. We limit computational expense of lifting and convolution by performing these operations over the truncated set $\{s_{-1},s_{0},s_{1}\}$. Model performance is reported in Table 3. The performance of our saturation-equivariant models out perform all other models, and the performance of our hue-equivariant model is on par with the baseline model.
@@ 5.4 CIFAR-10: Color shift in the wild
We evaluate the performance of our model in the presence of natural hue shifts on the CIFAR-10 [23] classification dataset. The dataset consists of 60k examples, 40k of which are used for training, 10k for validation, and 10k for testing. Model performance is reported in Table 4. The performance of our color-equivariant models are comparable to the baseline model.
@@ 6.1 Limitations and future work
While our method is equivariant to hue and saturation by design, it is not guaranteed to be equivariant to luminance. However, since luminance is computed as a linear combination of the red, green and blue channel values of an input image, if the preprocessing statistics (e.g., data mean) on the training set are valid for the testing set, luminance is approximately unchanged.
@@
7 Acknowledgements
We would like to thank members of the CAB Lab for valuable discussions and feedback, and the Princeton Department of Computer Science for providing compute resources. This work is partially supported by the Princeton School of Engineering and Applied Science.
@@ A.1 Hue-shift MNIST
We introduce the Hue-shift MNIST dataset to evaluate the performance of our equivariant architectures in the presence of global hue shift. We show select examples from the in-distribution and out-of distribution datasets in Figures 6(a) and 6(b).
@@ A.2 Hue-shift 3D Shapes
We introduce the Hue-shift 3D Shapes dataset to evaluate the performance of our equivariant architectures in the presence of local hue shift. We show select examples from the in-distribution and out-of distribution datasets in Figures 7(a)- 7(c).
@@ B.1 Hue-shift MNIST
We train our equivariant networks and the conventional architectures for 5 epochs with a batch size of 128. We optimize over a cross-entropy loss using the Adam optimizer [21] with $\beta_{1}=0.9$, and $\beta_{2}=0.999$. We use an initial learning rate of $10^{-3}$ for the $\mathbb{Z}^{2}$ network, and $10^{-4}$ for the hue-equivariant network. We train CEConv architectures using the hyperparameters reported in [29].
@@ B.2 Hue-shift 3D Shapes: Local hue equivariance
For this task we compare against the ResNet18 architecture proposed in [19], and the Z2CNN and ResNet44 architectures proposed in [7]. Our hue-equivariant architectures are designed with the same network structure as the baselines, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our hue-equivariant architectures, we perform hue group pooling to yield an hue-invariant representation.
Following [34], we train the ResNet architectures for 100k iterations with a batch size of 128. We optimize over a cross-entropy loss using SGD with a learning rate of $10^{-2}$. Images are down-sampled by a factor of 2 to train the CNN architectures. CNN architectures are trained as described in Section B.1.
@@ B.3 CIFAR-10: Color shift in the wild
For this task, we compare against the ResNet44 architecture proposed in [7]. Our hue-, saturation- and hue-saturation-equivariant architectures have the same number of layers as ResNet44, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our networks, we perform group pooling to yield group invariant representation.
We train our equivariant networks and the conventional architectures for 300 epochs, and a batch size of 128. We optimize over a cross-entropy loss using SGD with an initial learning rate of $10^{-1}$ and a cosine-annealing scheduler. We train CEConv architectures using the hyperparameters reported in [29].
@@ B.4 Camelyon17: Saturation shift in the wild
For this task, we compare against the ResNet50 architecture [19]. Our hue- and saturation-equivariant architectures have the same number of layers as ResNet50, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our hue- and saturation-equivariant architectures, we perform a group pooling to yield group invariant representation.
We train our equivariant networks and the conventional architectures for 10k iterations with a batch size of 32. We optimize over a cross-entropy loss using the Adam optimizer [21] with an initial learning rate of $10^{-2}$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$. We train CEConv architectures using the hyperparameters reported in [29].